OpenStudy (anonymous):

4(-Y+6x)=12 write this n a standard form?

Parth (parthkohli):

Just divide both sides by 4.

OpenStudy (anonymous):

so -4y+24x=12??

OpenStudy (anonymous):

wait divide??why?

Parth (parthkohli):

No!

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {4(-y + 6x) \over 4} = {12 \over 4} }\) \(\Large \color{MidnightBlue}{\Rightarrow 6x - y = 3 }\)

OpenStudy (anonymous):

Oo ok thank so thats a standard from?